# Perverse filtrations and Fourier transforms

### v2

@@ 5.3.2. Tautological classes and $P=C$
It was proven in [49] that the first $73r-$ 7 classes of cohomological degrees $\leq 2(r-2)$, i.e.,
@@
The bound $2(r-2)$ coincides with the bound obtained in Corollary 5.3. It is optimal in view of the free generation result [49]. However, this fact is crucial for our proof, but we do not have an easy geometric explanation of this coincidence. This bound is also optimal for the $P=C$ conjecture below; see [32, Remark 0.5].
@@ 6. $\bullet$
[…]
@@
Then $C\to W$ can be viewed as a family of curves in the linear system $|r\Sigma|$ with $\Sigma\subset S$ the 00-section. A universal sheaf $\overline{F}^{d}$ on $S\times\overline{J}^{d}_{C}$ of Section 5.2 provides a universal bundle

### v3

@@ 5.3.2. Tautological classes and $P=C$
It was proven in [49] that the first 3r73r-7 classes of cohomological degrees $\leq 2(r-2)$, i.e.,
@@
The bound 2(r2)2(r-2) coincides with the bound obtained in Corollary 5.3. It is optimal in view of the free generation result [49]. However, this fact is crucial for our proof, but we do not have an easy geometric explanation of this coincidence. This bound is also optimal for the $P=C$ conjecture below; see [32, Remark 0.5].
@@ 6. $\bullet$
(76) $M\xleftarrow{~~f~~}\widetilde{M}\xhookrightarrow{~~\widetilde{\iota}~~}M^{0}\xhookrightarrow{~~\iota_{0}~~}{M}^{\mathrm{par}}\xhookleftarrow{~~\iota~~}M^{\mathrm{parell}}\xrightarrow{~~q~~}\overline{J}^{d}_{C}.$
@@
Then $C\to W$ can be viewed as a family of curves in the linear system $|r\Sigma|$ with $\Sigma\subset S$ the 0-section. A universal sheaf $\overline{F}^{d}$ on $S\times\overline{J}^{d}_{C}$ of Section 5.2 provides a universal bundle